# hep-ph/***

### v1

@@ 1 Introduction
Many extensions of the SM offer a generation mechanism for neutrino masses. Among these are the well-known type-I [1, 2, 3, 4, 5], type-II [6, 7, 8, 9, 10, 11] and type-III seesaw mechanisms [12] which typically require new particles with masses much larger than the electroweak scale, if couplings are of order one. Alternatives such as the inverse seesaw (ISS) mechanism [13, 16, 14, 15] can reconcile new particles with masses in the TeV-range with not too small couplings, while the smallness of neutrino masses is due to a small breaking of lepton number (in the form of a Majorana mass term for the gauge singlets).
@@
In the current study, we analyse a scenario in which light neutrino masses are generated by the ISS mechanism with 3+3 gauge singlet fermions and the flavour symmetries $\Delta(3\,n^{2})$ and $\Delta(6\,n^{2})$ and CP are broken non-trivially.  [^1] We do not specify the breaking mechanism here. Concretely, the non-trivial flavour structure among the neutral states is encoded in the Dirac neutrino Yukawa matrix, while the mass matrices (Dirac and Majorana) of the gauge singlet fermions do neither break $G_{f}$ nor CP. This possibility has been called option 2 in [45]. We scrutinise analytically and numerically the Pontecorvo-Maki-Nakagawa-Sakata (PMNS) mixing matrix and different charged lepton flavour violating (cLFV) processes for Case 1) through Case 3 b.1). The main findings are that the effects of non-unitarity of the PMNS mixing matrix induced by the existence of the heavy sterile states have a more general form than those found in the scenario analysed in [45] (called option 1) and that signals of cLFV $\mu-e$ transitions can be sizeable in the studied parameter space, while those involving the tau lepton are more suppressed.
The remainder of the paper is organised as follows: in section 2 we outline the scenario, show the form of the different mass and coupling matrices, comment on the masses of the different neutral states, discuss the mixing matrices and the effects of the heavy sterile states on the light neutrinos. Section 3 contains information on lepton mixing including the characteristics of the mixing patterns of Case 1) through Case 3 b.1) and the effects of non-unitarity due to the mixing of the light neutrinos and the heavy sterile states. In section 4 we present analytic estimates as well as numerical results for the branching ratios (BRs) $\mathrm{BR}(\mu\to e\gamma)$ and $\mathrm{BR}(\mu\to 3\,e)$ and the $\mu-e$ conversion rate (CR) in aluminium, $\mathrm{CR}(\mu-e,\mathrm{Al})$, for examples of all cases, Case 1) through Case 3 b.1). We comment on the expected size of the signals of the cLFV tau lepton decays $\tau\to\mu\gamma$, $\tau\to e\gamma$, $\tau\to 3\,\mu$ and $\tau\to 3\,e$ in section 5 and summarise in section 6. In appendix A basics of the group theory of $\Delta(3\,n^{2})$ and $\Delta(6\,n^{2})$ can be found, while we collect the form of different relevant matrices in appendix B. Appendix C comprises a concise description of the numerical scan and supplementary plots displaying results of the studied $\mu-e$ transitions for Case 3 b.1) are given in appendix D.
@@ 2 Scenario
The group $G_{e}$ determines the structure of the charged lepton mass matrix $m_{\ell}$. We use a residual $Z_{3}$ symmetry that is a diagonal subgroup of the $Z_{3}$ symmetry, generated by the generator $a$ of the flavour group (which is represented by a diagonal matrix in the chosen basis, compare appendix A), and the additional $Z_{3}$ symmetry $Z_{3}^{\mathrm{(aux)}}$. Consequently, $m_{\ell}$ is diagonal and contains three free parameters corresponding to the masses of the electron, muon and tau lepton. We take these to be ordered canonically such that the contribution $U_{\ell}$ to lepton mixing is $U_{\ell}=\mathbb{1}$.
@@
where the second equality takes into account that $U_{\ell}=\mathbb{1}$ in the chosen basis. This non-unitarity is due to the mixing of the light neutrinos with the heavy states and it can be encoded in the hermitean matrix $\eta$,
@@
and the form of the hermitean matrix $\eta$ given by
@@
where the role of $K_{\nu}$ is to ensure that the light neutrino masses are positive semi-definite. [^4] We set $K_{\nu}$ to the identity matrix in analytic results, but take it properly into account in the numerical study. Interestingly, then the leading order values of the light neutrino masses $m_{f}$ are simply given by
@@ 3 Lepton mixing
In order to analyse the non-unitarity of the lepton mixing matrix, we extract the lepton mixing angles and CP invariants $J_{\mathrm{CP}}$, $I_{1}$ and $I_{2}$ from a three-by-three mixing matrix $U$ with elements $U_{\alpha i}$, $\alpha=e,\mu,\tau$ and $i=1,2,3$, as follows
@@
Before commenting on the results for the different cases, we discuss the generic effects using as form of the lepton mixing matrix either the one in eq. (35) or (43), depending on whether the matrix in square brackets in eq. (26) is diagonal or not. We define the deviations of the lepton mixing angles and CP invariants in the ISS framework as
$\Delta\sin^{2}\theta_{ij}=\frac{(\sin^{2}\theta_{ij})_{\mathrm{ISS}}-(\sin^{2}\theta_{ij})_{\mathrm{FCS}}}{(\sin^{2}\theta_{ij})_{\mathrm{FCS}}}\;,\;\;\Delta J_{\mathrm{CP}}=\frac{(J_{\mathrm{CP}})_{\mathrm{ISS}}-(J_{\mathrm{CP}})_{\mathrm{FCS}}}{(J_{\mathrm{CP}})_{\mathrm{FCS}}}\;,\;\;\Delta I_{i}=\frac{(I_{i})_{\mathrm{ISS}}-(I_{i})_{\mathrm{FCS}}}{(I_{i})_{\mathrm{FCS}}}\;,$ (45)
as long as the result obtained from the flavour and CP symmetries (FCS) is non-vanishing; this definition is identical to the one used in [45].
@@
where we refer with $U_{\alpha i}$ to the elements of the unitary matrix, i.e. the one obtained from flavour and CP symmetries. Clearly, in the limit in which $\theta_{L}$ and $\overline{\theta}_{L}$ coincide, $\Delta\theta=0$, we recover the formulae found in eq. (46). Given the more general form of the deviation it is possible that the non-unitary lepton mixing matrix leads to non-trivial CP violation, although flavour and CP symmetries predict certain CP phases to be trivial.
@@
As has been shown in [33], the choice of residual symmetries is described by one integer, called $s$, varying between 00 and $n-1$, which corresponds to the employed CP symmetry. Two of the three CP phases, $\beta$ and $\delta$, are trivial in this case, i.e. $I_{2}=0$ and $J_{\mathrm{CP}}=0$, while the third one, the Majorana phase $\alpha$, is determined by the choice of the CP symmetry, $|\sin\alpha|=|\sin 6\,\phi_{s}|$ with $\phi_{s}=\frac{\pi\,s}{n}$. [^6] The sign in the relation for the Majorana phase $\alpha$ depends on the explicit form of the matrix $K_{\nu}$. In contrast to this, the three lepton mixing angles are independent of this choice and only depend on the free angle $\theta_{L}$ (or in case the matrix combination in square brackets in eq. (26) is not diagonal this angle is given by $\overline{\theta}_{L}$). This angle is adjusted such that the reactor mixing angle is very close to its experimental best-fit value [49] and, consequently, the solar and atmospheric mixing angle are given by $\sin^{2}\theta_{12}\approx 0.341$ and $\sin^{2}\theta_{23}\approx 0.604\,(5)$ for light neutrino masses with NO (IO).
@@
From [33] it is known that the CP symmetry is fixed by two integer parameters, $s$ and $t$, ranging from 00 to $n-1$. Alternatively, one can define the combinations $u$ and $v$ as
@@
Using the results of [47] we know that for the parameters $m$ and $s$ both even or both odd the matrix combination in square brackets in eq. (26) is diagonal and the formulae in eq. (46) can be used. In contrast, for $m$ even and $s$ odd or vice versa this matrix combination is only diagonal for $y_{1}=0$ or $\sin 2\,\theta_{R}=0$. For $y_{1}=0$ we encounter the same situation as for Case 1) and Case 2), since light neutrino masses then follow strong NO and all deviations only depend on one Yukawa coupling $y_{f}$. If both parameters $m$ and $s$ are even or both are odd, setting $y_{3}=0$ permits us to suppress the deviations of the reactor and the atmospheric mixing angle, compare eq. (46). In case the matrix combination in square brackets in eq. (26) is not diagonal, it is diagonalised by a rotation in the $(12)$-plane. For this situation similar formulae as those in eqs. (47) and (48) can be derived.
@@
Like for Case 3 a), the integer parameters $m$ and $s$ with $10\leq m,s\leq n-$ 1 describe the residual $Z_{2}$ symmetry and the CP symmetry, respectively. From [33], we know that the requirement to accommodate the solar mixing angle well constrains the parameter $m$ to be close to $\frac{n}{2}$. The free angle, appearing in the lepton mixing matrix, has to be close to $\frac{\pi}{2}$ such that the reactor mixing angle can be fitted well. For $\frac{m}{n}=\frac{1}{2}$, usually two such values exist. Otherwise, it depends on the choice of $m$ and $s$ how many distinct values of the free angle admit a good fit to the experimental data of the lepton mixing angles. The atmospheric mixing angle also depends on the parameter $s$. Likewise, all CP phases are determined by the actual values of the ratios $\frac{m}{n}$ and $\frac{s}{n}$ as well as the free angle. More information, in particular about analytical estimates and numerical values of the CP phases, is given in [33, 45, 47]. Explicit choices of $m$ and $s$ are discussed in section 4.6.
@@ 4.1.1 Experimental inputs and constraints
The lepton mixing angles are constrained by the global fits found in [49] (version 5.3). [^9] In the numerical analysis we use version 5.2, but check that the fitted lepton mixing angles are also compatible with the results given in version 5.3. Since the experimental preference for a certain value of the Dirac phase $\delta$ is still rather weak, we do not include this information here, when fitting the angle $\theta_{L}$ in order to accommodate well the lepton mixing parameters.
@@
$m_{1}=m_{0}\;,\;\;m_{2}=\sqrt{m_{1}^{2}+\Delta m^{2}_{\mathrm{sol}}}\;,\;\;m_{3}=\sqrt{m_{1}^{2}+\Delta m^{2}_{\mathrm{atm}}}\;,$ (52)
@@ 4.1.2 Parameter space
In several of the following figures, see e.g. figs. 16 and 22, we only display results for values of $\theta_{R}$ that lie in the range $0\leq\theta_{R}\leq\frac{\pi}{4}$. Nevertheless, we have checked that we obtain the same results for values outside this range. The angle $\theta_{L}$, on the other hand, is determined by fitting the measured values of the lepton mixing angles as best as possible. Depending on the case, Case 1) through Case 3 b.1), we find one or two such values for $\theta_{L}$, see section 3. The couplings $y_{f}$ are fixed such that the neutrino masses $m_{i}$, $i=1,2,3$, are correctly reproduced, while still being perturbative, i.e. $y_{f}\leq\sqrt{8\,\pi}$ [59]. Indeed, in the presented scans these range in the interval $4\times 10^{-5}\lesssim y_{f}\lesssim 1.2$, independent of the ordering of the light neutrino masses. We choose the signs of $y_{f}$ to be positive. The scales $\mu_{0}$ and $M_{0}$ are varied in certain ranges. In particular, we take the lepton number violating parameter $\mu_{0}$ to be much smaller than the electroweak scale, as expected in the ISS framework, and vary it logarithmically as follows
@@
Heavy sterile states in such a mass range can have different observational imprints, see e.g. [60, 61, 62, 63, 64, 65, 66, 67, 68, 69]. In order to optimise the exploration of this range of $M_{0}$, we divide it into three intervals, i.e. a lower mass interval, $150\,\mbox{GeV}\leq M_{0}\leq 1\,\mbox{TeV}$, an intermediate mass interval, $1\,\mbox{TeV}\leq M_{0}\leq 5\,\mbox{TeV}$, and a higher mass interval, $5\,\mbox{TeV}\leq M_{0}\leq 10\,\mbox{TeV}$. In each of these intervals we vary $M_{0}$ uniformly and logarithmically. A concise description of the numerical scan is found in appendix C.

### v2

@@ 1 Introduction
Many extensions of the SM offer a generation mechanism for neutrino masses. Among these are the well-known type-I [1, 2, 3, 4, 5], type-II [6, 7, 8, 9, 10, 11] and type-III seesaw mechanisms [12] which typically require new particles with masses much larger than the electroweak scale, if couplings are of order one (and no symmetry is implemented, see e.g. [13, 14, 15]). Alternatives such as the inverse seesaw (ISS) mechanism [16, 19, 17, 18] can reconcile new particles with masses in the TeV-range with not too small couplings, while the smallness of neutrino masses is due to a small breaking of lepton number (in the form of a Majorana mass term for the gauge singlets).
@@
In the current study, we analyse a scenario in which light neutrino masses are generated by the ISS mechanism with 3+3 gauge singlet fermions and the flavour symmetries $\Delta(3\,n^{2})$ and $\Delta(6\,n^{2})$ and CP are broken non-trivially.  [^1] We do not specify the breaking mechanism here. Concretely, the non-trivial flavour structure among the neutral states is encoded in the Dirac neutrino Yukawa matrix, while the mass matrices (Dirac and Majorana) of the gauge singlet fermions do neither break $G_{f}$ nor CP. This possibility has been called option 2 in [48]. We scrutinise analytically and numerically the Pontecorvo-Maki-Nakagawa-Sakata (PMNS) mixing matrix and different charged lepton flavour violating (cLFV) processes for Case 1) through Case 3 b.1). The main findings are that the effects of non-unitarity of the PMNS mixing matrix induced by the existence of the heavy sterile states have a more general form than those found in the scenario analysed in [48] (called option 1) and that signals of $\mu-e$ transitions can be sizeable in the studied parameter space, while those involving the tau lepton are more suppressed.
The remainder of the paper is organised as follows: in section 2 we outline the scenario, show the form of the different mass and coupling matrices, comment on the masses of the different neutral states, discuss the mixing matrices and the effects of the heavy sterile states on the light neutrinos. Section 3 contains information on lepton mixing including the characteristics of the mixing patterns of Case 1) through Case 3 b.1) and the effects of non-unitarity due to the mixing of the light neutrinos and the heavy sterile states. In section 4 we present analytic estimates as well as numerical results for the branching ratios (BRs) $\mathrm{BR}(\mu\to e\gamma)$ and $\mathrm{BR}(\mu\to 3\,e)$ and the $\mu-e$ conversion rate (CR) in aluminium, $\mathrm{CR}(\mu-e,\mathrm{Al})$, for examples of all cases, Case 1) through Case 3 b.1). We comment on the expected size of the signals of the cLFV tau lepton decays $\tau\to\mu\gamma$, $\tau\to e\gamma$, $\tau\to 3\,\mu$ and $\tau\to 3\,e$ in section 5 as well as on the possible impact of experimental constraints from the study of Tritium beta ($\beta$) decay and searches for neutrinoless double beta ($0\nu\beta\beta$) decay in different nuclei in section 6. Eventually, we summarise in section 7. In appendix A basics of the group theory of $\Delta(3\,n^{2})$ and $\Delta(6\,n^{2})$ can be found, while we collect the form of different relevant matrices in appendix B. One can find information on the residual symmetries, the CP transformations and the mixing patterns for each case, Case 1) through Case 3 b.1), in appendix C. Appendix D comprises a concise description of the numerical scan and supplementary plots displaying results of the studied $\mu-e$ transitions for Case 2) and Case 3 b.1) are given in appendix E.
@@ 2 Scenario
The group $G_{e}$ determines the structure of the charged lepton mass matrix $m_{\ell}$. We use a residual $Z_{3}$ symmetry that is a diagonal subgroup of the $Z_{3}$ symmetry, generated by the generator $a$ of the flavour group (which is represented by a diagonal matrix in the chosen basis, compare appendix A), and the additional $Z_{3}$ symmetry $Z_{3}^{\mathrm{(aux)}}$. Consequently, $m_{\ell}$ is diagonal and contains three free parameters corresponding to the masses of the electron, muon and tau lepton. [^2] At the level of unbroken flavour and CP symmetries, the Yukawa couplings responsible for the charged lepton masses are expected to be of the form e.g. $\frac{y_{e}}{\Lambda}\,\overline{L}_{\alpha}\,H\,\phi_{e}\,\ell_{eR}$ with the phase of $y_{e}$ constrained by the imposed CP symmetry, $\Lambda$ being the cutoff scale of the theory and $\phi_{e}$ an SM gauge singlet that transforms as $({\bf 3},1)$ under the flavour symmetry and $Z_{3}^{\mathrm{(aux)}}$. If this scalar acquires a VEV that preserves the residual symmetry $G_{e}$, i.e. it has the form $\langle\phi_{e}\rangle=v_{e}\,\left(\begin{array}[]{c}1\\ 0\\ 0\end{array}\right)$ with $v_{e}$ in general complex, we obtain $\frac{y_{e}}{\Lambda}\,\overline{L}_{e}\,H\,v_{e}\,\ell_{eR}$ and after electroweak symmetry breaking the electron mass reads $m_{e}=\Big{|}\frac{y_{e}}{\Lambda}\,v_{e}\Big{|}\,\langle H\rangle$. The muon and tau lepton masses can be generated in an analogous way. For further information regarding this point, see, for example, [49] and appendix D of [50]. We take these to be ordered canonically such that the contribution $U_{\ell}$ to lepton mixing is $U_{\ell}=\mathbb{1}$.
@@
where the second equality takes into account that $U_{\ell}=\mathbb{1}$ in the chosen basis. This non-unitarity is due to the mixing of the light neutrinos with the heavy states and it can be encoded in the Hermitian matrix $\eta$,
@@
and the form of the Hermitian matrix $\eta$ is given by
@@
where the role of $K_{\nu}$ is to ensure that the light neutrino masses are positive semi-definite. [^5] We set $K_{\nu}$ to the identity matrix in analytic results, but take it properly into account in the numerical study. The explicit forms of mixing patterns can be found, for convenience, in appendix C. Interestingly, then the leading order values of the light neutrino masses $m_{f}$ are simply given by
@@ 3 Lepton mixing
In order to analyse the non-unitarity of the lepton mixing matrix, we extract the lepton mixing angles and CP invariants $J_{\mathrm{CP}}$, $I_{1}$ and $I_{2}$ from a three-by-three mixing matrix $U$ with elements $U_{\alpha i}$, $\alpha=e,\mu,\tau$ and $i=1,2,3$, as follows [^7] The formulae in eq. (44) correspond to those derived from a unitary mixing matrix that is parametrised by the three mixing angles $\theta_{12}$, $\theta_{13}$ and $\theta_{23}$ as well as the three CP phases $\delta$, $\alpha$ and $\beta$ in the following way $\begin{pmatrix}c_{12}c_{13}&s_{12}c_{13}&s_{13}e^{-i\delta}\\ -s_{12}c_{23}-c_{12}s_{23}s_{13}e^{i\delta}&c_{12}c_{23}-s_{12}s_{23}s_{13}e^{i\delta}&s_{23}c_{13}\\ s_{12}s_{23}-c_{12}c_{23}s_{13}e^{i\delta}&-c_{12}s_{23}-s_{12}c_{23}s_{13}e^{i\delta}&c_{23}c_{13}\end{pmatrix}\,\left(\begin{array}[]{ccc}1&0&0\\ 0&e^{i\alpha/2}&0\\ 0&0&e^{i(\beta/2+\delta)}\end{array}\right)$ with $s_{ij}=\sin\theta_{ij}$ and $c_{ij}=\cos\theta_{ij}$.
@@
Before commenting on the results for the different cases, we discuss the generic effects using as form of the lepton mixing matrix either the one in eq. (35) or (43), depending on whether the matrix in square brackets in eq. (26) is diagonal or not. We define the (relative) deviations of the lepton mixing angles and CP invariants in the ISS framework as
$\Delta\sin^{2}\theta_{ij}=\frac{(\sin^{2}\theta_{ij})_{\mathrm{ISS}}-(\sin^{2}\theta_{ij})_{\mathrm{FCS}}}{(\sin^{2}\theta_{ij})_{\mathrm{FCS}}}\;,\;\;\Delta J_{\mathrm{CP}}=\frac{(J_{\mathrm{CP}})_{\mathrm{ISS}}-(J_{\mathrm{CP}})_{\mathrm{FCS}}}{(J_{\mathrm{CP}})_{\mathrm{FCS}}}\;,\;\;\Delta I_{i}=\frac{(I_{i})_{\mathrm{ISS}}-(I_{i})_{\mathrm{FCS}}}{(I_{i})_{\mathrm{FCS}}}\;.$ (45)
Here, the quantities labelled with “ISS” correspond to those extracted from (in general non-unitary) lepton mixing matrices which have a form as found in e.g. eq. (35), while the ones carrying the label “FCS” (flavour and CP symmetries) are obtained from (unitary) lepton mixing matrices whose forms are determined by the mismatch of the residual symmetries $G_{e}$ and $G_{\nu}$ (without assuming a specific model for generating Majorana masses for the light neutrinos), compare the lepton mixing patterns found in e.g. [36] (see also appendix C). The definition in eq. (45) is only valid, if the relevant quantity with “FCS” is non-vanishing. We note that this definition is identical to the one used in [48].
@@
where we refer here with $U_{\alpha i}$ to the elements of the unitary matrix, i.e. the one obtained from flavour and CP symmetries. Clearly, in the limit in which $\theta_{L}$ and $\overline{\theta}_{L}$ coincide, $\Delta\theta=0$, we recover the formulae found in eq. (46). Given the more general form of the deviation it is possible that the non-unitary lepton mixing matrix leads to non-trivial CP violation, although flavour and CP symmetries predict certain CP phases to be trivial.
@@
As has been shown in [36] (compare also appendix C), the choice of residual symmetries is described by one integer, called $s$, varying between 00 and $n-1$, which corresponds to the employed CP symmetry. Two of the three CP phases, $\beta$ and $\delta$, are trivial in this case, i.e. $I_{2}=0$ and $J_{\mathrm{CP}}=0$, while the third one, the Majorana phase $\alpha$, is determined by the choice of the CP symmetry, $|\sin\alpha|=|\sin 6\,\phi_{s}|$ with $\phi_{s}=\frac{\pi\,s}{n}$. [^8] The sign in the relation for the Majorana phase $\alpha$ depends on the explicit form of the matrix $K_{\nu}$. In contrast to this, the three lepton mixing angles are independent of this choice and only depend on the free angle $\theta_{L}$ (or in case the matrix combination in square brackets in eq. (26) is not diagonal this angle is given by $\overline{\theta}_{L}$). This angle is adjusted such that the reactor mixing angle is very close to its experimental best-fit value [54] and, consequently, the solar and atmospheric mixing angle are given by $\sin^{2}\theta_{12}\approx 0.341$ and $\sin^{2}\theta_{23}\approx 0.604\,(5)$ for light neutrino masses with NO (IO).
@@
From [36] it is known that the CP symmetry is fixed by two integer parameters, $s$ and $t$, ranging from 00 to $n-1$; see appendix C for its explicit form. Alternatively, one can define the combinations $u$ and $v$ as
@@
Using the results of [52] we know that for the parameters $m$ and $s$ both even or both odd the matrix combination in square brackets in eq. (26) is diagonal and the formulae in eq. (46) can be used. In contrast, for $m$ even and $s$ odd or vice versa this matrix combination is only diagonal for $y_{1}=0$ or $\sin 2\,\theta_{R}=0$. For $y_{1}=0$ we encounter the same situation as for Case 1) and Case 2), since light neutrino masses then follow strong NO and each deviation only depends on one Yukawa coupling $y_{f}$. If both parameters $m$ and $s$ are even or both are odd, setting $y_{3}=0$ permits us to suppress the deviations of the reactor and the atmospheric mixing angle, compare eq. (46). In case the matrix combination in square brackets in eq. (26) is not diagonal, it is diagonalised by a rotation in the $(12)$-plane. For this situation similar formulae as those in eqs. (47) and (48) can be derived.
@@
Like for Case 3 a), the integer parameters $m$ and $s$ with $10\leq m,s\leq n-$ 1 describe the residual $Z_{2}$ symmetry and the CP symmetry, respectively. From [36], we know that the requirement to accommodate the solar mixing angle well constrains the parameter $m$ to be close to $\frac{n}{2}$; consult appendix C for the explicit form of the mixing matrix. The free angle, appearing in the lepton mixing matrix, has to be close to $\frac{\pi}{2}$ such that the reactor mixing angle can be fitted well. For $\frac{m}{n}=\frac{1}{2}$, usually two such values exist. Otherwise, it depends on the choice of $m$ and $s$ how many distinct values of the free angle admit a good fit to the experimental data of the lepton mixing angles. The atmospheric mixing angle also depends on the parameter $s$. Likewise, all CP phases are determined by the actual values of the ratios $\frac{m}{n}$ and $\frac{s}{n}$ as well as the free angle. More information, in particular about analytical estimates and numerical values of the CP phases, is given in [36, 48, 52]. Explicit choices of $m$ and $s$ are discussed in section 4.6.
@@ 4.1.1 Experimental inputs and constraints
The lepton mixing angles are constrained by the global fits found in [54]. [^11] In the numerical analysis we use version 5.2, but check that the fitted lepton mixing angles are also compatible with the results given in version 5.3. Since the experimental preference for a certain value of the Dirac phase $\delta$ is still rather weak, we do not include this information here, when fitting the angle $\theta_{L}$ in order to accommodate well the lepton mixing parameters. For all displayed points of the numerical scans, e.g. for Case 1) in fig. 3, the lepton mixing angles agree at the $3\,\sigma$ level or better with the global fit results [54].
@@
$m_{1}=m_{0}\;,\;\;m_{2}=\sqrt{m_{0}^{2}+\Delta m^{2}_{\mathrm{sol}}}\;,\;\;m_{3}=\sqrt{m_{0}^{2}+\Delta m^{2}_{\mathrm{atm}}}\;,$ (52)
@@ 4.1.2 Parameter space
In several of the following figures, see e.g. figs. 16 and 30, we only display results for values of $\theta_{R}$ that lie in the range $0\leq\theta_{R}\leq\frac{\pi}{4}$. Nevertheless, we have checked that we obtain the same results for values outside this range. The angle $\theta_{L}$, on the other hand, is determined by fitting the measured values of the lepton mixing angles as best as possible. Depending on the case, Case 1) through Case 3 b.1), we find one or two such values for $\theta_{L}$, see section 3. The couplings $y_{f}$ are fixed such that the neutrino masses $m_{i}$, $i=1,2,3$, are correctly reproduced, while still being perturbative, i.e. $y_{f}\leq\sqrt{8\,\pi}$ [66]. Indeed, in the presented scans these range in the interval $4\times 10^{-5}\lesssim y_{f}\lesssim 1.2$, independent of the ordering of the light neutrino masses. We choose the signs of $y_{f}$ to be positive. The scales $\mu_{0}$ and $M_{0}$ are varied in certain ranges. In particular, we take the lepton number violating parameter $\mu_{0}$ to be much smaller than the electroweak scale, as expected in the ISS framework, and vary it logarithmically as follows
@@
Heavy sterile states in such a mass range can have different observational imprints, see e.g. [67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81]. In order to optimise the exploration of this range of $M_{0}$, we divide it into three intervals, i.e. a lower mass interval, $150\,\mbox{GeV}\leq M_{0}\leq 1\,\mbox{TeV}$, an intermediate mass interval, $1\,\mbox{TeV}\leq M_{0}\leq 5\,\mbox{TeV}$, and a higher mass interval, $5\,\mbox{TeV}\leq M_{0}\leq 10\,\mbox{TeV}$. In each of these intervals we vary $M_{0}$ uniformly and logarithmically. A concise description of the numerical scan is found in appendix D.